OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol. plz help

OpenStudy (anonymous):

I'm not sure about the formula you gave, but this is how I would normally approach a percent problem... So you have Chris, and he wants a $100 coat for some reason. Since he's an employee, he will receive a 10% discount on top of the 15% sale. To find the price of the coat on sale (without employee discount) you would take 100 X .15 I don't currently have a calculator, so you'll have to work that out. Lol To find the price including the employee discount, you would take the amount you just found on your calculator X .10 Now that you have both of those numbers, add them together, and then subtract them from the price of the coat. :)

OpenStudy (anonymous):

C = 0.765x + 0.06(0.765x) X = 100 C = 0.765(100) + 0.06(0.765*100) C = 76.50 + 0.06(76.50) C = 76.50 + 4.59 C = 81.09 (sale price of coat)